kjallday:

The employees of a company have different hobbies. 14 men who like playing golf 6 women who like playing golf 2 men who like running 18 women who like running Which statement is correct? A. For every man who likes running, 7 men like playing golf. B. For every woman who likes running, 7 women like playing golf. C. For every man who likes running, 3 men like playing golf. D. For every woman who likes running, 9 women like playing golf.

kittybasil:

Let's expand on the parameters for each choice. A. "For every man who likes running, 7 men like playing golf." We know 2 men like running and 14 men like playing golf. So there is a ratio of 2 running to 14 golf, or\[2:14\rightarrow\frac{2}{14}\]We can simplify this ratio. Both the top and bottom are divisible by 2:\[\frac{2\div2}{14\div2}=\frac{1}{7}\]So, here our ratio is 1 running to 7 golf. Which is exactly the same as the statement you see.

kittybasil:

B. "For every woman who likes running, 7 women like playing golf." We know 18 women like running and 6 women like playing golf. So there is a ratio of 18 women to 6 golf, or\[18:6\rightarrow\frac{18}{6}\]We can simplify this ratio, because both top and bottom are divisible by 6:\[\frac{18\div6}{6\div6}=\frac{3}{1}\]This gives you a ratio of 3 running to 1 golf. That isn't what the choice gives you, so this is wrong.

kittybasil:

C. "For every man who likes running, 3 men like playing golf." See my first reply (below)... the ratio does not match, so this is also incorrect.

1 attachment
kittybasil:

D. "For every woman who likes running, 9 women like playing golf." See my second reply (below)... the ratio does not match, so this is also incorrect.

1 attachment
kittybasil:

Based on the previous information, can you figure out which answer choice is correct?